OpenStudy (anonymous):

The sum of two numbers is 61, and their difference is sixteen more than the smaller number. What are the two numbers?

OpenStudy (anonymous):

Begin by listing out your givens.

OpenStudy (asnaseer):

call one number x and the other one y, where x > y. then we can write: x + y = 61 x - y = 16y and solve to find x and y.

OpenStudy (anonymous):

I got this part but Im lost at what to do next

OpenStudy (anonymous):

I'm not understanding why you have 2 equations. Let x = smaller number Let x + 16 = larger number

OpenStudy (asnaseer):

the second part of the question states: "...their difference is sixteen more than the smaller number" which translates to: x - y = 16y

OpenStudy (anonymous):

Would it not be x - y = 16 + y then?

OpenStudy (asnaseer):

hmmm.... maybe you are right here - I took "16 more" to mean "16 times more". I think your interpretation is the correct one.